OpenStudy (anonymous):

Find the solutions to the equation (check all that apply) 20x2-26x+8=0. a.x=1/5 b.x=1/2 c.x=4/5 d.x=2/3 e.x=1/3 f.x=3/5

OpenStudy (anonymous):

b. x=1/2 c. x=4/5